Line two is perpendicular to the line 2x-4y=20. What is the slope of the line?
Answer by blwinbbbles(106)

First put the equation into linear form of y = mx + b
so:
2x - 4y = 20 subtract 2x from both sides
-4y = -2x + 20 now divide by -4
y = 1/2(x) - 5 so you now know the slope is 1/2
a line perpendicular has an inverse negative slope:
so if the slope were -1/2 the perpendicular line would have a slope of 2
In our case the slope is 1/2 so the P line would have a slope of -2
Hope this helps